Session: May 2001 (session 2)
Wavelength: 18. cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (A.Maccaferri at Mon May 28 04:49:05 METDST 2001) Success
all ok

Torun (K.Borkowski at Mon May 28 09:26:30 METDST 2001) Success
No known problems.

Onsala (A. Polatidis/J. Conway at Mon May 28 10:02:30 METDST 2001) Success

Westerbork (Tony Foley at Mon May 28 11:35:54 METDST 2001) Minor failures
Started 18minutes late because of a problem with the delay settings of the tied array. Otherwise OK

Effelsberg (A. Kraus at Mon May 28 15:58:05 METDST 2001) Minor failures
Frequently HPE. Several times we were late on source - insufficient slewing time. No other known problems.

Noto (P. Leto at Wed May 30 13:13:34 METDST 2001) Minor failures
23:05:12UT - 23:18:12UT data lost for tape change problem

Jodrell_Bank (Alastair Gunn at Wed May 30 17:55:59 METDST 2001) Success
No known problems.